OpenStudy (anonymous):

find x and y intercept -5x+9y=-45 please show work

OpenStudy (anonymous):

it's easier if you re-write your equation as: 9y=5x-45 y=5/9 x-5 x-intercept means y=0: x=9 (9,0) y-intercept means x=0: y=-5 (0,-5)

OpenStudy (anonymous):

When x=0, 9y=-45, y=-5 When y=0, -5x=-45, x=-9